# Non-perturbative Origin of Electroweak Scale via Higgs-portal: Dyson-Schwinger in Conformally Invariant Scalar Sector

### v1

@@ I Introduction
Among several possibilities including supersymmetry and extra-dimensional theories one popular and quite elegant solution to this problem is to hypothesize scale invariance as a symmetry of the fundamental action, such that all scales we observe in Nature be generated dynamically beyond classical level. Coleman and Weinberg [17] proposed such a scenario and showed that the SM gauge symmetry breaking could be radiatively triggered via quantum corrections. However, from the observational point-of-view this mechanism fails within the Standard Model to generate the correct Higgs mass (the Electroweak (EW) Scale) as it predicts $m_{Z,W}>m_{H}$, where $m_{Z,W}$ are the SM Z and W gauge boson masses while $m_{H}$ is the SM Higgs boson mass [17, 21]. Inspite of this faliure this has remained the direction of BSM model building and several extensions of the SM has been explored extensively in the literature [1, 17, 86, 18, 20, 19], where this radiative EW symmetry breaking (EWSB) mechanism is successful in terms of observations and often makes very concrete and testable predictions. When non-minimal coupling to gravity is introduced, such scenarios can provide naturally flat inflaton potentials [77, 74, 84, 86, 72, 71, 7, 89] and stable particle dark matter candidates [62, 75, 72, 73, 76, 6]. The scenarios also lead to very strong first-order phase transitions and hence the possibility of high amplitude detectable gravitational wave (GW) signals in upcoming detectors [70, 82, 69, 5, 83, 11, 81, 60]. Consequently, scale invariant scenarios offer an interesting direction of model-building for solving the hierarchy problem in the Standard Model of particle physics [24, 2, 21, 62, 22, 3, 65, 86, 18, 72, 23, 73]. See Refs. [63, 4, 66, 67, 68, 12, 10] for other studies of conformal invariance and dimensional transmutation of energy scales along similar lines [85, 13, 52, 45, 59, 43, 55, 57].
@@ III.2 Gap equations
[…]
@@
We obtain the plot in Fig.3 for the inequality (49). The red curve is well below the Higgs boson mass for a large set of values of $\lambda_{\phi}$ as required and so, there exists a meaningful range of parameters for which the scenario is fully consistent.

### v2

@@ I Introduction
Among several possibilities including supersymmetry and extra-dimensional theories one popular and quite elegant solution to this problem is to hypothesize scale invariance as a symmetry of the fundamental action, such that all scales we observe in Nature be generated dynamically beyond classical level. Coleman and Weinberg [18] proposed such a scenario and showed that the SM gauge symmetry breaking could be radiatively triggered via quantum corrections. However, from the observational point-of-view this mechanism fails within the Standard Model to generate the correct Higgs mass (the Electroweak (EW) Scale) as it predicts $m_{Z,W}>m_{H}$, where $m_{Z,W}$ are the SM Z and W gauge boson masses while $m_{H}$ is the SM Higgs boson mass [18, 22]. In spite of this failure this has remained the direction of BSM model building and several extensions of the SM has been explored extensively in the literature [1, 18, 87, 19, 21, 20], where this radiative EW symmetry breaking (EWSB) mechanism is successful in terms of observations and often makes very concrete and testable predictions. When non-minimal coupling to gravity is introduced, such scenarios can provide naturally flat inflaton potentials [78, 75, 85, 87, 73, 72, 7, 90] and stable particle dark matter candidates [63, 76, 73, 74, 77, 6]. The scenarios also lead to very strong first-order phase transitions and hence the possibility of high amplitude detectable gravitational wave (GW) signals in upcoming detectors [71, 83, 70, 5, 84, 11, 82, 61]. Consequently, scale invariant scenarios offer an interesting direction of model-building for solving the hierarchy problem in the Standard Model of particle physics [25, 2, 22, 63, 23, 3, 66, 87, 19, 73, 24, 74]. See Refs. [64, 4, 67, 68, 69, 12, 10] for other studies of conformal invariance and dimensional transmutation of energy scales along similar lines [86, 13, 53, 46, 60, 44, 56, 58].
@@ III SM Higgs mass from hidden sector mass gap
Before to go through the full model, it would be helpful to understand the decoupled case with $\beta=0$. One is left with a quartic scalar theory. This case has been extensively studied by us in [40, 16] and two kinds of solutions are expected. The quartic model admits an exact solutions in terms of Jacobi elliptical functions for eq.(II) and all higher order correlation functions are expressed through them that is the hallmark of a Gaussian solution. On the other hand, as proven in [16], there is also the constant solution that arises from quantum fluctuations as in eq.(II) the term $G_{2}(0)$, when properly evaluated and regularized, yields a mass term. This term has the right sign to give rise to the Higgs mechanism as currently understood for the Standard Model. Thus, it is up to nature to decide what solution applies. For simplicity reasons, the Higgs mechanism appears the most economical one and seems the one observed in experiments. Indeed, the other solution implies an infinite tower of massive excitations that were not observed so far [^3] A comparison with the Coleman-Weinberg effective potential [18] could only be possible in the approximate limit of a small coupling while, in our case, we have an exact solution that holds at any value of the coupling itself..
@@ III.2 Gap equations
$m_{n}=(2n+1)\frac{\pi}{2K(i)}\left(\sqrt[4]{\frac{\lambda_{\phi}}{2}}\mu+\frac{m_{\phi}^{2}}{2^{\frac{3}{2}}\sqrt[4]{\lambda_{\phi}}\mu}\right).$ (48)
@@
We obtain the plot in Fig.3 for the inequality (49). The red curve is well below the Higgs boson mass for a large set of values of $\lambda_{\phi}$ as required and so, there exists a meaningful range of parameters for which the scenario is fully consistent as we get the correctly observed EW Higgs mass observed at LHC, being $\lambda_{\phi}$ fixed (within the uncertainty of Higgs mass data), for $\mu\sim 1.5\ {\rm TeV}$ and the condition (49) granted.
@@ V Acknowledgements
Data sharing
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest
The authors declare no conflict of interest.